# Low Resource Reconstruction Attacks Through Benign Prompts

### v2

@@ Abstract
In this work, we devise a new attack that requires low resources, assumes little to no access to the actual training set, and identifies, seemingly, benign prompts that lead to potentially-risky image reconstruction. This highlights the risk that images might even be reconstructed by an uninformed user and unintentionally. For example, we identified that, with regard to one existing model, the prompt “blue Unisex T-Shirt” can generate the face of a real-life human model. Our method builds on an intuition from previous works which leverages domain knowledge and identifies a fundamental vulnerability that stems from the use of scraped data from e-commerce platforms, where templated layouts and images are tied to pattern-like prompts. [^1] Our attack implementation is publicly available at https://github.com/TheSolY/lr-tmi.
@@ 1 Introduction
With the increasing use of generative and foundation models, and the increasing use of public and potentially private data for training, a range of new risks and challenges has evolved within the field of Machine Learning. Today, there is growing concern that the use of data may lead to breaches of copyright and violations of individual privacy. While privacy is typically associated with sensitive or non-public data, even public data raises important concerns [34]. Individuals who have consented to share their photos publicly may still expect control over how their data is used, particularly to prevent it from being presented in unintended or inappropriate contexts.
Source Generated Prompt
Refer to caption Refer to caption "Galaxy Area Rug"
(source: GearFrost)
Refer to caption Refer to caption "Floral iPhone Case"
(source: eBay)
Refer to caption Refer to caption "Abstract Art Shower Curtain"
(source: Amazon)
Figure 1: Examples of template memorized images reconstructed through our attack on SD 1.4.
@@
In turn, researchers are now investigating how data is memorized within generative and foundation models, and whether data can be extracted in ways unintended. Several works [33, 13, 36] demonstrated that attacks can be designed to extract, blatantly and verbatim, data that appeared in the training sets. These existing attacks tend to rely on access to the training data, some on substantial computational resources, and they search for specific prompts, from training data that, potentially, trigger the extraction. In that sense, these attacks simulate a malicious adversary that explicitly aims to extract such data. However, an important concern is the potential for unintentional image conjuring, where a user issues a benign prompt that might inadvertently trigger the same phenomenon.
@@
Thus, towards a better understanding of the potential risks, we develop in this work a followup attack. Our objective is to construct simple and benign prompts that call for generic objects such as t-shirt, round metal wall art, beach towel etc., without too-specific details. We applied our attack on previously attacked model, and use these prompts to extract images that present elements and objects whose source can be traced to other websites. Perhaps the most disturbing feature of our attack is that real-life models, whose images appeared in such websites, are also extracted by these, so-called innocent prompts, as seen in Fig. 4. In comparison, previous attacks could extract real images of real people but typically relied on intentional prompts, and the scope was to generate the exact image. For instance, Carlini et al. [13] showed that prompting with “Ann Graham Lotz” could yield a verbatim copy of an existing image. The objective was not to show that her image will be extracted (as the prompt requests that) but to investigate the copying of the existing photo. In contrast, our attack shows that even unintentional prompts can generate images of real individuals without explicitly requesting them. This behavior raises distinct concerns around privacy, and individuals’ rights. Particularly the right not to have their likeness used for modeling purposes without their consent.
@@
Discussion and Limitations
Extracting training data presents a major challenge, raising serious concerns around privacy and potential copyright infringement. Our results focus on the phenomenon of unintentional extraction of memorized data. Our attack demonstrates that image memorization in diffusion models can be exploited with minimal resources and without access to the training dataset, posing a more widespread privacy and copyright risk than previously understood. Because our attack relies on seemingly innocuous prompts (e.g., "Abstract Art Unisex T-Shirt") we reveal that users may unintentionally generate memorized images of real individuals or copyrighted content. These types of attacks amplify the risks associated with inadvertent copying and highlight the troubling possibility that even seemingly benign users may inadvertently trigger such leaks.
The vulnerability we exploit arises from the template-structure of scraped e-commerce data, where uniform layouts across product categories facilitate memorization and make leakage more detectable. While engineers already attempt to clean datasets of obvious duplicates, our attack suggests that duplication should be interpreted in a more conservative sense, that also accounts for repeated patterns and templates. Most of our attacks were conducted on the SD 1.4 version, but even in the most recent models such as Stable Diffusion v3.5, Flux Schnell, and Midjourney v6, we could still obtain some of the results from version 1.4, as presented in Section 4.2. We could also trace some memorization of previously reconstructed images, as presented in Appendix B. It is possible that methods of decoupling between text and image that were presented in the training process of version 3.5 significantly strengthened the model against the threat of our attack, but it did not completely mitigate it. The examples for successful extraction of template-memorized images from version 3.5 which are shown in Fig. 6 hint that it is not completely immune to our attack. We leave it to future study to further improve our attack, as well as to develop more principled methods to protect models from similar attacks.
As we discuss, the ability to extract memorized data poses various potential risks. In this work we focused solely on extracting data, or variants thereof, that is already accessible. Specifically, all images of identifiable individuals that are presented in this work are limited to cases where the original image is publicly available, easily accessible on the internet and the alterations are limited to benign visual modifications—such as changes in clothing color or background.
@@
In recent years several reconstruction attacks for foundation models have been designed, including attacks targeted on language models [12, 25, 24] and, more aligned with out work, image reconstruction [33, 13, 36]. We elaborate further on these attacks and how they relate to our work in Section 2.
We are specifically concerned with attacks that generate data from prompts, or black-box (to various extent) models that through the “standard” use of the model induce types of image conjuring. This is distinct from various attacks that reconstruct training data by analyzing and probing the model [19, 39, 17]. These works highlight the memorization of training data by large learning models. Such memorization is necessary for reconstruction attacks to even be possible. There is increasing theoretical support that such memorization may be unavoidable [26, 10, 35, 16], and reconstruction attacks complement these works by showing that such memorization has implications. This leads to the open problem of mitigating risks and prevent data reconstruction even when the model inherently retains information from its training data.
Given that data is memorized, an important aspect that arises is the question of originality, or how to make sure that the output of the model is original and not mere interpolation of existing data. This leads to an interesting theoretical question – to formalize and model what constitute original data [15, 30], and to an imminent practical question – to identify and regulate non-originality [20, 18, 14]. Our attack generates results in a middle ground between blatant, verbatim copying and non-copying that lacks originality, where the system replicates elements without directly copying. Such types of interpolations have been observed before [9, 32]. It is possible that such interpolations can be mitigated and regulated through appropriate credit attribution. The question of how to properly attribute credit, though, is an ongoing research question [27].
@@ 2 Background and Existing Attacks
To select the candidate captions they relied on full duplicate of image-text pairs. The candidates were selected as the highest scoring 2 million image-text pairs from LAION2B on the duplication metric presented in [37]. The 2 million candidate include only fully duplicated text-image pairs, and does not include images that were duplicated in a partial fixed region, along with a partial text, but for which no full duplication existed. Out of the 2M candidates, 30k were selected based on their score in the white-box attack presented at [36], i.e. those for which the most noise had been removed at the first denoising step of Stable Diffusion V1. Due to the method of identifying duplicates, they identified a phenomena of template memorization, where only a certain object in the image is being duplicated. They are also the first to postulate that such images might be traced to e-commerce sites, a fact that we further validate and helped us in devising our attack that does not require to harvest duplications from the training dataset. Therefore, we expand on the phenomena of template memorization, showing that one could use a natural English text including a short Collocation rather than a highly specific text, and that such Collocations can extract several image templates rather than a specific one. We further rely on e-commerce websites to select candidate prompts and collocations, releasing us from searching and using full duplicates, thus expanding the search to image-text pairs that were duplicated only in part.
@@ 3.1 Data Collection
For each of these sites, we look at the snapshot of the site from March 2021, since the latest Common Crawl version to enter LAION 5B was April 2021. We then scraped a list of collocations from categories that are sold on these websites such as “unisex t-shirts", “athletic shoes" etc. Overall we reached a list of around $108$ collocations. In e-commerce websites, images for each category are generated through different templates. Therefore, in order to generate diversity, that will allow us to identify duplicates, we generated, for each collocation, a set of prompts by adding a description-visual pattern to each. The descriptions were: "Galaxy", "Floral, "Abstract Art" , "I Heart ML", "blue", "red". This process leads to a set of prompts, pattern+collocation phrases, such as “Floral Unisex t-shirts". For each such prompt we generate $25-50$ different images (distinct by their seeds from 00 to $49$).
@@
Only for the sake of comparison with previous work, we also extracted list of categories that were related to sources that previous work also identified. These categories allow us to further analyze our attack and enable comparisons. In more detail, we took the prompts already identified by Webster [36] as what they denote "templates verbatim". These are images memorized up to a fixed mask, or "retrieval verbatim"222these are identified by generated images in a near duplicate of a training image but the caption of the training image is different than the prompt used to generate the generated image.. Similarly we extracted collocations from Somepalli et al. [32], Hintersdorf et al. [21] that we identified as product descriptions (rather than news titles etc.). From these, we manually isolated prompted collocations as the 1-3 words long phrase describing a product. For example, in the list of prompts memorized by Stable Diffusion, published by Webster [36], appears the prompt "Outshak Hand-Knotted Beige/Blue Area Rug by Shalom Brothers" out of which we extracted the collocation "Area Rug". After collecting a list of candidate expressions, we generated many images similar to our previous method where we add to the collocation a description pattern.
@@ 3.2 Searching near-duplicates
We then run clique search based on the similarity of the predicted fixed region, where adjacent images are pairs with cosine similarity of above 0.95 on the CLIP embedding. Clique of at least size 2 were suspected as template-memorized, and were filtered to identify near-duplicates. In addition, using CLIP instead of $\ell_{2}$ distance, included images with perturbations which are minor visual variations between images (see Section 5.2), that are otherwise very similar in the same clique. We also searched for near-duplicates through manual visual inspection to complete the results missing from the limitation of the segmentation models.
“X Area Rug” (“rug”) Refer to caption
@@
“X Shower Curtain” (“curtain”) Refer to caption
“X High-Tops Sneakers” (“Right-shoe”) Refer to caption
@@ 3.3 Tracing the Source Image
As we could narrow down the search of the source to handful number of websites, some images were recognized even without duplications, such as the generated image shown in first row of Fig. 4. For this reason, we then continued to visually inspect the generated images comparing them to the original images in the origin websites, under their product categories, leading to exposing of more such memorized images.
@@ 4 Results
All of our experiments were conducted on a single RTX A6000 Machine. We primarily focused on Stable Diffusion version 1.4 from the official checkpoint accessible at HuggingFace [4], which is the model where previous attacks were introduced [13, 36, 33]. Results from more recent models are discussed and reported in Section 4.2.
In Figure 2 we demonstrate examples of four cliques we identified in four different collocations. Since, we generate the cliques using CLIP embedding and not $\ell_{2}$ distance, notice that in the curtain example cliques may include slightly zoomed-in versions of the same image. On the other hand, some of the duplicates (e.g. the rugs) can differ only in the color of the rug (that is being segmented and masked for clique identification), while others may involve differing patterns (for example, the wall tapestries).
@@ 4.1 Images with identified source
For many of the duplicated images, we could locate a source image from the web. Some examples are depicted in Fig. 3. Noteworthy, even for some images that we did not identify them as duplicated, we could identify a source image by inspecting e-commerce sites from which we’ve taken the categories, for example the image on the first row of Fig. 4.
Category Abstract Art Galaxy Floral I Heart ML Source
@@
T-Shirt Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
T-Shirt Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Examples across product categories.
In Fig. 8, we show reconstructed images matched LAION sources and their captions, even though the prompts used were generic and the captions themselves. This highlights the risk of unintentionally reproducing training data.
@@
Arguably, the most concerning type of image recall are those which potentially contain a real human model. Several of the duplicates we identified indeed contain real humans that we could find their source, see Fig. 4.
As discussed, previous works also managed to extract real-humans [13]. However this was generally done by prompts that actively request that person’s image, and the focus was on copyright and verbatim copy. Here, the person is extracted by a prompt that, seemingly, does not contain their name, but instead the prompt contains a collocation associated with a product that the model had presented in the image. For example, mundane items such as "T-Shirt". Even in cases where the face was deformed (as often happens in SD) or the picture is only of the torso, it might still contain identifying markers such as tattoos as well as other specific visual attributes such as haircut and pose. This is exemplified in Fig. 9 (left columns).
Source & Generated Image Prompt Used & Source
Refer to caption “blue Unisex T-Shirt”
@@
Refer to caption “Abstract Art A-Line Dress”
@@
Refer to caption “Abstract Art T-Shirt”
@@
Figure 4: Examples of generated people whose source images were identified (left: source, right: generated). The first row shows a person identified by name [11].
4.2 Attacks on State-Of-The-Art and other models
While we focused out efforts, and designed our attack on Stable Diffusion 1.4 (SD 1.4), we could also extract template-memorized images from other popular generative models. Primarily, by using the same collocations identified in SD 1.4. In Fig. 5 we exhibit further images extracted from DeepFloyd IF-XL-I-v1.0 [1] and Midjourney V4 [3]. Both models were released around the time of SD 1.4, and in Fig. 6 we demonstrate further extractions from state-of-the-art (SOTA) models.
  • DeepFloyd IF-XL-I-v1.0: although the generated images exhibited lower contrast, which made clique search impractical, we successfully identified template-memorized images through visual inspection. In addition, we discovered new templates—absent from SD 1.4—using Google Lens searches on manually flagged generations.
    For each prompt and seed, we collected the four default outputs returned by Midjourney, specifying version 4 with the argument --v 4. Despite limited prompt and seed control, we successfully extracted template-like generations from Midjourney. Our results included both direct template replications from SD 1.4 and interpolated reconstructions (see Section 5.1).
    DeepFloyd
    (a) Example results from our attack on DeepFloyd. For each pair, the source image (left) and the generated image (right) are shown.
    Midjourney v4
    (b) Example results from our attack on Midjourney v4. For each pair, the source image (left) and the generated image (right) are shown.
    Figure 5: Reconstructed images from models from the same generation as SD 1.4
    For Stable diffusion and flux Schnell we used the Hugging Face checkpoint in the same manner we’ve conducted our main experiments on SD v1.4, while for Midjourney we used the Discord interface as was detailed previously. We successfully demonstrated some memorization, as seen in Fig. 6, though to a lesser extent than in older models and versions.
    SD 3.5 introduced partial mitigation techniques targeting image-text coupling vulnerabilities; however, these measures were not specifically designed to address the template-style coupling exploited by our attack, which relies on concise prompts derived from external sources rather than training data. Our results indicate that SD 3.5 exhibits increased resilience but remains partially vulnerable to template memorization. Representative examples are provided in Figure 6
    MidJourney V6.1, as discussed, poses additional challenges. However, we applied the same methodology used for V4. While the generated images were typically perturbed (Section 5.2), targeting previously vulnerable benign prompts still yielded images with notable resemblance to the source. For instance, as shown in Fig. 6, the prompt “Abstract Art Essential T-Shirt" (which previously produced near-verbatim copies) continues to generate closely resembling images.
Stable Diffusion v3.5 Flux-Schnell v1.0 MidJourney v6.1
Source Generated Source Generated Source Generated

### v3

@@ Abstract
In this work, we devise a new attack that requires low resources, assumes little to no access to the actual training set, and identifies, seemingly, benign prompts that lead to potentially-risky image reconstruction. We further show that image reconstructions may occur even unintentionally and by users without specific knowledge. For example, we observed that for one existing model, the prompt “blue Unisex T-Shirt” generates the face of a real individual. Moreover, by combining the identified vulnerabilities with real-world prompt data, we were able to find prompts that reproduce memorized elements.
Our method builds on an intuition from previous works which leverages domain knowledge and identifies a fundamental vulnerability that stems from the use of scraped data from e-commerce platforms, where templated layouts and images are tied to pattern-like prompts.
The code for our attack is publicly available at https://github.com/TheSolY/lr-tmi.
@@ 1 Introduction
With the rising popularity of generative and foundation models, large amounts of public and sometimes private data are being used for training. This widespread use of data raises pressing concerns about privacy and copyright, as models may inadvertently reproduce protected content or reveal sensitive information. As such models become increasingly embedded in everyday applications, addressing these risks has become an urgent challenge for the field. In addition to concerns regarding copyright, the use of public data also raises important questions about consent and context [34]. Individuals who have consented to use their photos in a specific setting may not anticipate, nor approve of, their data being repurposed in unrelated or inappropriate contexts.
Source Generated
Refer to caption Refer to caption
Figure 1: An image that we generated on a SD 1.4 with the prompt "Walter White wall mural" that was extracted from the database of real-world used prompts DiffusionDB [8] (right). On its left a corresponding source image the contains elements that were copied. See Section˜4.5 for further details.
In turn, researchers are now investigating how data is memorized within generative and foundational models, and whether data can be extracted in ways unintended. Several works [31, 5, 37] demonstrated that attacks can be designed to extract, blatantly and verbatim, data that appeared in the training sets.
@@
These existing attacks typically rely on access to the training data, substantial computational resources, and the use of specific prompts drawn from that data to trigger extraction. In that sense, they simulate a malicious adversary explicitly attempting to recover training content. However, a critical concern lies in the potential for unintentional image reconstruction, where a user issues a benign prompt that inadvertently reproduces training data.
@@
Thus, towards better understanding of the potential risks, we develop in this work a followup attack. Our objective is to construct simple and benign prompts that call for generic objects such as t-shirt, round metal wall art, beach towel etc., without too-specific details. We applied our attack on previously attacked model, and use these prompts to extract images that present elements and objects whose source can be traced to other websites. Perhaps the most disturbing feature of our attack is that real-life models, whose images appeared in such websites, are also extracted by these, so-called innocent prompts, as seen in Figure˜6. In comparison, previous attacks could extract real images of real people but typically relied on highly specific prompts, typically refering to the individual by name, and the scope was to generate the exact image. For instance, Carlini et al. [5] showed that prompting with “Ann Graham Lotz” could yield a verbatim copy of an existing image. The objective was not to show that her image will be extracted (as the prompt requests that) but to investigate the copying of the existing photo. In contrast, our attack shows that even unintentional prompts can generate images of real individuals without explicitly requesting them (see Figure˜6). Moreover, using the results from our attack, as well as existing data-bases for real-world used prompts, we could allocate apparently harmless prompts that users indeed entered and can generate images containing copied elements (see LABEL:fig:gen-vs-source). These behaviors raise distinct concerns around privacy, and individuals’ rights. Particularly the right not to have their likeness used for modeling purposes without their consent, or usage of copyrighted material.
@@
In recent years several reconstruction attacks for foundation models have been designed, including attacks targeted on language models [4, 19, 18] and, more aligned with out work, image reconstruction [31, 5, 37]. We elaborate further on these attacks and how they relate to our work in Section˜2.
We are specifically concerned with attacks that generate data from prompts, or black-box (to various extent) models that through the “standard” use of the model induce types of image conjuring. This is distinct from various attacks that reconstruct training data by analyzing and probing the model [14, 39, 12]. These works highlight the memorization of training data by large learning models. Such memorization is necessary for reconstruction attacks to even be possible. There is increasing theoretical support that such memorization may be unavoidable [21, 2, 35, 10], and reconstruction attacks complement these works by showing that such memorization has implications. This leads to the open question of how to mitigate risks and prevent data reconstruction even when the model inherently retains information from its training data.
Given that data is memorized, an important aspect that arises is the question of originality, or how to make sure that the output of the model is original and not mere interpolation of existing data. This leads to a very interesting theoretical question of what constitute original data [9, 27], and to an imminent practical question as to how to identify and regulate non-originality [15, 13, 6]. Our attack generates results in a middle ground between blatant, verbatim copying and non-copying that lacks originality, where the system replicates elements without directly copying. While similar interpolations have been observed previously [1, 30], our method is specifically designed to reliably produce such reconstructions at scale. It is possible that such interpolations can be mitigated and regulated through appropriate credit attribution. The question of how to properly attribute credit, though, is an ongoing research question [20].
@@ 2 Background and Existing Attacks
Another issue is looking for cliques at least 10 images, we’ve seen empirically that some generated images that we could identify as extracted training images by the similarity to a specific training image, were generated less that 10 times, and even only once. We challenge the implicit assumption of one-to-one memorization, where a specific text is associated with a specific image, causing the specific image to return many times, by showing how some collocations are associated with different image templates, such that the generated images are a mixture of those template along with novel generations.
@@
To select the candidate captions they relied on full duplicate of image-text pairs. The candidates were selected as the highest scoring 2 million image-text pairs from LAION2B on the duplication metric presented in [36]. The 2 million candidate include only fully duplicated text-image pairs, and does not include images that were duplicated in a partial fixed region, along with a partial text, but for which no full duplication existed. Out of the 2M candidates, 30k were selected based on their score in the white-box attack presented at [37], i.e. those for which the most noise had been removed at the first denoising step of Stable Diffusion V1. Due to the method of identifying duplicates, they identified a phenomena of template memorization, where only a certain object in the image is being duplicated. They are also the first to postulate that such images might be traced to e-commerce sites, a fact that we further validate and helped us in devising our attack that does not require to harvest duplications from the training dataset. Therefore, we expand on the phenomena of template memorization, showing that one could use a natural English text including a short key phrase which we call a collocation rather than a highly specific text, and that such collocations can extract several image templates rather than a specific one. We further rely on e-commerce websites to select candidate prompts and collocations, releasing us from searching and using full duplicates, thus expanding the search to image-text pairs that were duplicated only in part.
@@ 3.1 Data Collection
For each of these sites, we look at the snapshot of the site from March 2021, since the latest Common Crawl version to enter LAION 5B was April 2021. We then scraped a list of collocations from categories that are sold on these websites such as "unisex t-shirts", “athletic shoes" etc. Overall we reached a list of around $108$ collocations. In e-commerce websites, images for each category are generated through different templates. Therefore, in order to generate diversity, that will allow us to identify duplicates, we generated, for each collocation, a set of prompts by adding a description-visual pattern to each. The descriptions were: "Galaxy", "Floral, "Abstract Art" , "I Heart ML", "blue", "red". This process leads to a set of prompts, pattern+collocation phrases, such as "Floral Unisex t-shirts". For each such prompt we generate 255025-50 different images (distinct by their seeds from 0 to $49$).
Source Generated Prompt
Refer to caption Refer to caption "Galaxy Area Rug"
(source: GearFrost)
Refer to caption Refer to caption "Floral iPhone Case"
(source: eBay)
Refer to caption Refer to caption "Abstract Art Shower Curtain"
(source: Amazon)
Figure 2: Examples of template-memorized images reconstructed through our attack on SD 1.4. For each pair, the source image (left), the generated image (center), and the corresponding text prompt (right) are shown.
@@
Only for the sake of comparison with previous work, we also extracted list of categories that were related to sources that previous work also identified. These categories allow us to further analyze our attack and enable comparisons. In more detail, we took the prompts already identified by Webster [37] as what they denote "templates verbatim". These are images memorized up to a fixed mask, or "retrieval verbatim" [^1] These are identified by generated images in a near duplicate of a training image but the caption of the training image is different than the prompt used to generate the generated image.. Similarly we extracted collocations from Somepalli et al. [30], Hintersdorf et al. [16] that we identified as product descriptions (rather than news titles etc.). From these, we manually isolated prompted collocations as the 1-3 words long phrase describing a product. For example, in the list of prompts memorized by Stable Diffusion, published by Webster [37], appears the prompt "Outshak Hand-Knotted Beige/Blue Area Rug by Shalom Brothers" out of which we extracted the collocation "Area Rug". After collecting a list of candidate expressions, we generated many images similar to our previous method where we add to the collocation a description pattern.
@@ 3.2 Searching near-duplicates
We then run clique search based on the similarity of the predicted fixed region, where adjacent images are pairs with cosine similarity of above 0.95 on the CLIP embedding. Clique of at least size 2 were suspected as template-memorized, and were filtered to identify near-duplicates. In addition, using CLIP instead of $\ell_{2}$ distance, included images with perturbations which are minor visual variations between images (see Section˜5.2), that are otherwise very similar in the same clique. The method described in this section was clearly limited by the accuracy of the segmentation models, returning only partial results. We leave it for future research to run similar experiments with better segmentation models. We also searched for near-duplicates through manual visual inspection to complete the results missing from the limitation of the segmentation models.
"X Area Rug" ("rug") Refer to caption
@@
"X Shower Curtain" ("curtain") Refer to caption
"X High-Tops Sneakers" ("Right-shoe") Refer to caption
@@ 3.3 Tracing the Source Image
As we could narrow down the search of the source to handful number of websites, some images were recognized even without duplications, such as the generated image shown in first row of Figure˜6. For this reason, we then continued to visually inspect the generated images comparing them to the original images in the origin websites, under their product categories, leading to exposing of more such memorized images.
@@ 4 Results
Figure 4: Number of collocations tested vs the number of identified image templates, mean of 5 permutations.
All of our experiments were conducted on a single RTX A6000 Machine and, excluding those in Figure˜9, used Stable Diffusion version 1.4 from the official checkpoint accessible at HuggingFace [32], which is the model where previous attacks were introduced [5, 37, 31]. Results for more recent models are discussed and reported in Section˜4.3.
The 112 collocations that we tested produced 63 image templates that we could identify, the dispersion of images is relatively uniform, as is summarized in Figure˜4. In particular, our method of choosing collocations provides constant marginal returns. Overall, we generated 11,40011{,}400 images—comparable to the 9,0009{,}000 produced by [30], but orders of magnitude fewer than the $175$ million generated in [5]. Beyond that, whereas prior works focused on individual reconstructions - each counted as a single text–image pair - we identified templates, encompassing a vast range of possible variations enabled by the editability attribute highlighted in this As shown in LABEL:tab:prompt_comparison, compared with previous approaches, our method more accurately pinpoints prompts to reproduce copied images. Finally, we validate our manual procedure for identifying copied images through a user study, summarized in Section˜4.4..
The full list of template groups that we identified, collocations and their associated image template, are provided in listed at Appendix˜A.
@@ 4.1 Images with identified source
In Figure˜3, we show examples of four cliques identified across different collocations. Since the cliques are formed using CLIP embeddings rather than raw $\ell_{2}$ distance, the grouped images may exhibit minor variations such as zoom, color, or pattern changes.
For many duplicated images, we were able to locate their source counterparts on the web, as shown in Figure˜5. Interestingly, even for some images not automatically identified as duplicates, we could still trace their origins by inspecting e-commerce sites from which our categories were derived—for example, the image in the first row of Figure˜6.
Category Abstract Galaxy Floral I Heart ML Source
@@
T-Shirt (1) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
T-Shirt (2) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Examples across product categories, showing different prompt themes (columns) and their corresponding real source image (last column).
In Figure˜12, we show two cases where the source images and their original captions could be identified in LAION, as they were borrowed from prior works with dataset access. The reconstruction prompts we used did not include these captions but rather generic descriptions, underscoring the risk of unintentional image conjuring.
@@
Arguably, the most concerning type of image recall are those which potentially contain a real human model. Several of the duplicates we identified indeed contain real humans that we could find their source, see Figure˜6.
Generated Image & Source Prompt Used & Source
Refer to caption "Abstract Art Unisex T-Shirt"
@@
Refer to caption "Abstract Art Graphic T-Shirt Dress"
@@
Refer to caption “Abstract Art Essential T-Shirt”
@@
Figure 6: Examples of generated people that we could identify in a source (left: generated image, right: source). The first row contains a person identified by name [3]. Attack conducted on SD 1.4.
As discussed, previous works also managed to extract real-humans [5]. However this was generally done by prompts that actively request that person’s image, and the focus was on copyright and verbatim copy. Here, the person is extracted by a prompt that, seemingly, does not contain their name, but instead the prompt contains a collocation associated with a product that the model had presented in the image. For example, mundane items such as "T-Shirt". This is both concerning as these models were compensated for generating exactly such images, and as an uninformed user could generate such images under the impression that the human image is synthetic.
Even in cases where the face was deformed (as often happens in SD) or the picture is only of the torso, it might still contain identifying markers such as tattoos as well as other specific visual attributes such as haircut and pose. See Figure˜15 (left columns) for further details.